OpenStudy (anonymous):

Find the volume of' the solid formed by revolving the region bounded by the graph of f(x)=-x^(2)+2x and the x-axis about the x-axis...

OpenStudy (amistre64):

you will want your x intercepts as your limits of integration

OpenStudy (anonymous):

I want to check my ans...my ans is \[\frac{ -304pi }{ 15 }\]

OpenStudy (amistre64):



OpenStudy (amistre64):

when doing rotations about an axis, you most likely want to use the sum of areas of a circle: pi r^2, from a to b; where r = f(x)\[\int_{a}^{b}\pi[f(x)]^2~dx\]

OpenStudy (amistre64):

and volume is never negative

OpenStudy (amistre64):

since your x intercepts are 0 and 2\[\int_{0}^{2}\pi[-x^2+2x]^2~dx\] \[\int_{0}^{2}\pi(x^4-4x^3+4x^2)~dx\]

OpenStudy (amistre64):

if i did it right on the paper .... 14/3 pi

OpenStudy (amistre64):

i messed up the adding :)

OpenStudy (amistre64):

16/15 pi

OpenStudy (anonymous):

I forget a step...haha the correct ans is 16/15 pi ~ thank you, you answer my question very detail... I know something new too~^^